# Exact Simulation of Longitudinal Data from Marginal Structural Models

### v2

@@ Appendix C Competing risk
Parameter Scenarios
Table6 summarizes the parameter values for treatment effect size and confounding scenarios:
@@
Runtime details
Table 7 reports the runtime required to simulate data under the setup for the large causal effect with medium confounding scenario. The simulations were carried out on a local laptop using a single score without parallelization. Generating data for $n$ up to 10,000 units takes only a few seconds, while even for a very large sample size ($n=50,\!000$) and 11 time steps ($K=10$), the runtime remains under one minute. This computational efficiency is crucial for tasks such as method evaluation and sample size calculations in clinical trial design, where simulations must be repeated across various parameter values. For example, determining the sample size might require testing 10 different parameter combinations (such as effect sizes, prevalence of drop-in treatments, or adherence levels) across four sample sizes, each bootstrapped 200 times to estimate statistical power. This requires $10\times 4\times 200=8,\!000$ simulations. Using our method, suppose $n=10,000$ and $K=10$, the total runtime is approximately 13 hours on a single core, which can be reduced to about 2 hours using an 8-core laptop or under 10 minutes on a 50-core CPU through parallelization.
In contrast, simulation methods requiring computationally intensive operations, such as Monte Carlo approximations (see, for example, Seaman and Keogh, 2024) or numerical integration, are often too slow to support such large-scale simulations. This highlights a key advantage of our proposed method: its ability to handle extensive simulation demands efficiently, making it well-suited for practical applications.
$n$ $K=5$ $K=10$ 1,000 0.27 0.79 2,000 0.45 1.68 5,000 1.69 2.49 10,000 2.70 6.09 50,000 16.50 45.49
Table 7: Runtime (in seconds) to simulate a dataset of size $n$, averaged over 20 repeats.
Weight calculation
To fit the Cox marginal model, we employ stabilized inverse probability of treatment weighting, with weights for subject $i$ at time step $k$ calculated as:
@@
The denominator in the weight formula represents the correctly specified propensity score model under the assumed data-generating process, while the numerator stabilizes the weights. To avoid unintended confounding, the baseline covariates included in the numerator must also be adjusted for in the marginal structural model (Cole and Hernán, 2008). In accordance with the model in (15), we only include $C$ in the numerator for stabilization.

### v3

@@ Appendix E Seaman and Keogh (2024) simulations
We target a Cox marginal structural model given by:
$\lambda(\widetilde{y}_{k+1}\,|\,do(\bar{a}_{k}),c,y_{k}=0)=\lambda_{0}\exp\left(\beta_{C}\,c+\beta_{A}\,a_{k}+\beta_{CA}\,c\,a_{k}\right),$ (23)
where $\lambda_{0}=\exp(-2)$ is the baseline hazard. We set $\beta_{C}=-0.5$ and vary the values of $(\beta_{A},\beta_{CA})$ to construct scenarios with different strengths of treatment: a strong effect scenario $(-0.5,0.3)$ and a weak effect scenario $(0.2,0.1)$. This parameterization assumes a constant hazard rate between follow-up visits, influenced by the treatment status at the most recent visit. This implies that given survival at follow-up time $k$, the incremental survival time $\widetilde{Y}_{k+1}$ follows an exponential distribution with the rate equal to the hazard above:
$\widetilde{Y}_{k+1}\sim\operatorname{Exp}(\lambda_{0}\exp\left(\beta_{C}\,c+\beta_{A}\,a_{k}+\beta_{CA}\,c\,a_{k}\right)).$
If $\widetilde{Y}_{k+1}<1$ then $Y_{k+1}=1$, and 00 otherwise. We set the total number of follow-ups $K$ to 6. We set the copulas describing the association between $L_{j}$ ($j=0,\ldots,k$) and $Y_{k+1}$ to be Student t with parameter $\rho_{L}$ and 5 degrees of freedom, although each pairwise dependence could, in principle, be distinct. Similarly, we model the copulas capturing the dependence between $Y_{k+1}$ and each $W_{j}$ ($j=0,\ldots,k$), given $L_{k}$, as Gaussian with correlation $\rho_{W}$, though again we note that different copulas could be specified for each time point. We vary the values of $\rho_{L}$ and $\rho_{W}$ to construct scenarios with different strengths of confounding. Table 7 summarizes the combinations of treatment effect size and confounding levels considered.
@@
We simulate data for sample size $n=\{1000,2000,5000\}$.
We contrast two approaches for fitting the Cox marginal structural model specified in (23): stabilized inverse-probability-weighting and unweighted estimation. The first employs stabilized inverse probability of treatment weighting, with weights for subject $i$ at time step $k$ calculated as:
@@
Table 8 summarizes the bias of estimates from 500 replications, compared to corresponding Monte Carlo standard errors as a measure of uncertainty. As the sample size increases, generally, the biases of estimates for $\beta_{C}$, $\beta_{A}$ and $\beta_{AC}$ obtained with inverse probability weights converge to zero, with their absolute values substantially smaller than the corresponding Monte Carlo standard errors. In contrast, the unweighted estimator exhibits considerable bias, particularly for the estimate of the treatment effect, $\beta_{A}$, which persists even at $n=5000$. The magnitude of this bias far exceeds the Monte Carlo standard errors. The unweighted estimates are most biased under the high level of confounding by $L$ and $W$, since this is not adjusted for.
Level of confounding Effect size $n$ $\beta_{C}$ $\beta_{A}$ $\beta_{AC}$ IPTW Unweighted IPTW Unweighted IPTW Unweighted High $1000$ $0.00(2.18)$ $-0.74(1.46)$ $-0.16(2.26)$ $2.75(1.88)$ $0.21(2.99)$ $0.22(2.20)$ $\beta_{A}=-0.5$ $2000$ $-0.07(1.50)$ $-0.74(1.00)$ $-0.13(1.56)$ $2.71(1.25)$ $0.24(2.06)$ $0.24(1.51)$ $5000$ $-0.04(1.13)$ $-0.78(0.67)$ $-0.02(1.03)$ $2.89(0.80)$ $0.04(1.46)$ $0.11(1.02)$ $1000$ $0.09(2.26)$ $-0.51(1.60)$ $0.27(1.92)$ $3.00(1.56)$ $-0.16(2.76)$ $0.07(2.02)$ $\beta_{A}=0.2$ $2000$ $0.04(1.64)$ $-0.63(1.15)$ $0.07(1.54)$ $2.84(1.15)$ $-0.02(2.04)$ $0.21(1.47)$ $5000$ $-0.04(1.07)$ $-0.73(0.70)$ $0.03(0.91)$ $2.81(0.68)$ $0.03(1.26)$ $0.31(0.88)$ Medium $1000$ $-0.13(2.12)$ $-0.77(1.48)$ $0.08(2.40)$ $2.66(1.78)$ $0.00(3.12)$ $0.20(2.23)$ $\beta_{A}=-0.5$ $2000$ $-0.01(1.60)$ $-0.79(0.97)$ $0.03(1.72)$ $2.53(1.25)$ $0.10(2.29)$ $0.39(1.52)$ $5000$ $0.00(1.04)$ $-0.74(0.60)$ $-0.04(1.09)$ $2.58(0.79)$ $0.10(1.52)$ $0.30(0.96)$ $1000$ $0.01(2.34)$ $-0.70(1.56)$ $0.28(1.97)$ $2.63(1.54)$ $-0.11(2.90)$ $0.30(2.05)$ $\beta_{A}=0.2$ $2000$ $0.01(1.73)$ $-0.67(1.04)$ $0.06(1.38)$ $2.50(1.08)$ $0.01(2.02)$ $0.35(1.37)$ $5000$ $-0.02(1.12)$ $-0.73(0.68)$ $0.05(1.03)$ $2.46(0.66)$ $0.05(1.41)$ $0.41(0.86)$ Low $1000$ $-0.09(2.04)$ $-0.11(1.41)$ $-0.05(2.40)$ $0.64(1.75)$ $0.12(3.12)$ $-0.12(2.25)$ $\beta_{A}=-0.5$ $2000$ $0.02(1.57)$ $-0.12(0.97)$ $0.02(1.70)$ $0.63(1.22)$ $0.09(2.26)$ $0.05(1.50)$ $5000$ $-0.06(1.07)$ $-0.12(0.65)$ $-0.01(1.11)$ $0.70(0.77)$ $0.08(1.47)$ $-0.07(0.94)$ $1000$ $0.14(2.21)$ $0.04(1.44)$ $0.27(2.09)$ $0.72(1.54)$ $-0.16(2.85)$ $-0.14(1.92)$ $\beta_{A}=0.2$ $2000$ $-0.02(1.75)$ $-0.08(1.05)$ $0.01(1.62)$ $0.58(1.11)$ $0.12(2.23)$ $-0.01(1.41)$ $5000$ $0.11(1.03)$ $-0.07(0.60)$ $0.10(0.89)$ $0.58(0.64)$ $-0.11(1.32)$ $-0.02(0.81)$
Table 9 presents coverage rates of the 95% confidence intervals. Confidence intervals for the weighted estimator, computed using both sandwich variance estimation and bootstrap resampling, achieve coverage close to the nominal level, with slight undercoverage. In contrast, the sandwich-based confidence intervals for the unweighted estimator perform poorly, especially for the treatment effect estimate, $\beta_{A}$.
Level of confounding Effect Size $n$ $\beta_{C}$ $\beta_{A}$ $\beta_{AC}$ W-S W-B UW W-S W-B UW W-S W-B UW High 1000 93.5% 94.3% 92.5% 93.8% 93.5% 60.3% 94.3% 93.8% 95.0 % $\beta_{A}=-0.5$ 2000 94.8% 95.3% 89.8% 95.3% 93.3% 39.0% 95.3% 93.8% 95.5 % 5000 92.3% 92.0% 75.5% 93.8% 94.5% 5.8% 94.0% 92.8% 93.5 % 1000 93.5% 94.3% 93.0% 95.5% 95.5% 49.3% 94.5% 94.3% 93.0 % $\beta_{A}=0.2$ 2000 92.5% 93.0% 90.0% 93.8% 92.5% 26.5% 94.0% 93.5% 92.8 % 5000 93.5% 94.0% 82.0% 94.8% 93.5% 1.0% 94.8% 94.3% 94.0 % Medium 1000 94.8% 94.8% 91.3% 94.3% 95.0% 64.0% 93.5% 93.3% 95.8 % $\beta_{A}=-0.5$ 2000 94.3% 94.3% 88.8% 93.8% 93.8% 44.3% 95.0% 94.0% 95.8 % 5000 97.5% 96.8% 83.0% 95.0% 94.8% 8.8% 94.3% 93.5% 92.8 % 1000 93.3% 93.3% 91.8% 93.8% 93.5% 58.3% 92.0% 93.8% 94.8 % $\beta_{A}=0.2$ 2000 93.5% 94.8% 91.0% 96.5% 96.8% 36.3% 94.8% 95.8% 93.5 % 5000 96.0% 94.5% 79.8% 94.5% 93.8% 6.5% 96.0% 95.0% 91.8 % Low 1000 93.5% 93.8% 94.0% 93.5% 93.3% 94.0% 93.5% 94.0% 94.5 % $\beta_{A}=-0.5$ 2000 94.8% 95.3% 94.0% 95.8% 95.3% 92.5% 94.0% 94.5% 95.8 % 5000 93.8% 94.3% 92.3% 95.3% 95.0% 86.0% 95.3% 93.8% 95.0 % 1000 93.0% 94.8% 94.8% 93.3% 94.5% 92.8% 95.3% 95.8% 94.5 % $\beta_{A}=0.2$ 2000 92.0% 92.3% 94.0% 92.5% 93.8% 90.8% 92.3% 94.0% 94.0 % 5000 96.5% 96.2% 96.0% 95.2% 95.2% 88.5% 95.0% 94.5% 96.5 %